OpenStudy (subbie):

Can a matrix of 5x5 matrices be in the subspace of 5x5 matrices? For a matrix of 5x5 matrices: 1. Why are invertible matrices not in the Subspace of 5x5 matrices? Eg. Invertible matrix A + invertible matrix B = matrix C that may not be invertible. So matrix C can not be in the subspace? $A.B$= matrix is not invertible as there will never be a zero matrix in A or B? So $A.B$ will never be in the subspace? 2. Why are singular matrices not in the Subspace of 5x5 matrices? Singular matrix D + singular matrix E = singular matrix F that may not be singular. So matrix F can not be in the subsp

OpenStudy (loser66):

Down the problem into smaller dimension to see what is going on. \[A= \left[\begin{matrix}1&0\\1&1\end{matrix}\right]\rightarrow A^{-1}=\left[\begin{matrix}1&0\\-1&1\end{matrix}\right]\] \[B=\left[\begin{matrix}1&0\\1&-1\end{matrix}\right]\rightarrow B^{-1}=\left[\begin{matrix}1&0\\1&-1\end{matrix}\right]\] Hence both A, B are invertible. But. \[C=A+B=\left[\begin{matrix}2&0\\2&0\end{matrix}\right]\] is not invertible.

OpenStudy (subbie):

Thank you for explaining =)